# NeRF-CA: Dynamic Reconstruction of X-ray Coronary Angiography with Extremely Sparse-views

### v1

@@ IV NeRF-CA
Figure 2: Overview of the NeRF-CA method (a). The loss functions are (1) Photometric Loss $\mathcal{L}_{p}$, (2) Static vs. Dynamic Factorization Loss $\mathcal{L}_{b}$, (3) Dynamic Entropy Loss $\mathcal{L}_{e}$, and (5) Dynamic Occlusion Loss $\mathcal{L}_{o}$. We also depict the (b) training input and (c) validation output, in yellow and pink, respectively.
In this section, we describe our method, NeRF-CA, which allows for dynamic Coronary Angiography reconstruction with extremely sparse-views. Figure 2 (a) shows an overview of our method. We categorize the components of our method according to their function in reconstructing the scene. Specifically, we utilize dynamic reconstruction components, shown in green in Figure 2 (a), and sparse-view reconstruction components, shown in purple in Figure 2 (a). Overall, the X-ray scene is expressed as an absorption-only model with attenuation coefficients $\sigma$, as will be further described in Section IV-B. Figure 2 (b) and (c) depict the training input and validation output of our model in yellow and pink, respectively. We will utilize these colors throughout the paper to distinguish the training and validation views.
The base of our method is decoupling the scene in a static and dynamic scene. The static model represents the background, and the dynamic model represents the dynamic coronary artery structure. The static and dynamic models each predict their attenuation coefficient $\sigma_{s}$ and $\sigma_{d}$, respectively. These coefficients are composited to obtain the scene from which the predicted pixel intensities are calculated. Figure 2 (c) depicts these outputs of our methods, specifically, the static, dynamic, and composite outputs for one validation view. Whereas the static model solely relies on the positionally-encoded 3D coordinate as input, the dynamic models also receive a per-cardiac phase latent code $\tau_{i}$. This latent code represents a cardiac phase along the cardiac cycle derived from the simultaneously recorded ECG signal. More details on the decomposition of the scene can be found in Section IV-C.
By decomposing the scene, we can individually impose regularizers for the static and dynamic scenes, allowing the final dynamic scene to model the coronary artery. In Section IV-A, we provide the overall objective of our method introducing these regularizers. Sections IV-D and IV-E describe the regularizers for dynamic and sparse-view reconstruction, respectively. The positional encoding of our method improves both dynamic and sparse-view reconstruction, indicated by the two colors in Figure 2. We also perform weighted pixel sampling to guide the model to distinguish the sparse blood vessel structure from the background. We perform this sampling with a blood vessel probability map, representing the pixel probability of having a blood vessel. This probability map is shown in Figure 2 (b) and detailed in Section IV-D. We provide the final combined loss function $\mathcal{L}_{f}$ in Section IV-F.
@@ IV-A Overall Objective
Our overall objective consists of two parts: dynamic and sparse-view reconstruction. Our contribution lies in adapting and combining these dynamic and sparse-view reconstruction strategies. To demonstrate these strategies, we provide a running example throughout the explanation of our method. Figure 2 (c) shows examples of the predicted coronary angiogram views and the ground truth views.
For dynamic reconstruction, we decouple the static background from the dynamic coronary artery structure, as detailed in Section IV-D. We combine standard static and dynamic decomposition with a weighted pixel sampling strategy and two regularizers to decouple the coronary artery from the background accurately. Weighted pixel sampling is introduced to separate the sparse blood vessel structure from the background on the image level. The Static vs. Dynamic Factorization regularizer $\mathcal{L}_{b}$ is enforced to separate the static and dynamic scene on a 3D point level [11]. On top of that, we enforce entropy $\mathcal{L}_{e}$ in the decoupled dynamic scene to achieve the sparse structure of the blood vessel [12].
For sparse-view reconstruction, we reconstruct from an extremely sparse amount of angiogram sequences. The sparse-view reconstruction components, consisting of windowed positional encoding and dynamic occlusion loss, will be detailed in Section IV-E. The windowed positional encoding is a proven strategy to enforce smoothness while combating degenerate high-frequency solution [8]. We enforce this strategy on both the static and dynamic scenes. The occlusion loss $\mathcal{L}_{o}$ punishes high occlusion areas close to the camera, an expected faulty behavior when applying NeRFs for sparse-view reconstruction [8]. We demonstrate that imposing such a loss can benefit our sparse dynamic scene.
@@ IV-B Neural Radiance Field for X-ray
X-ray scenes can be described by the absorption-only model, assuming the Beer-Lambert Law [20]. Similarly to other works applying NeRFs for X-ray scenes (e.g., [31]), we approximate the scene by $F_{\Theta}(\boldsymbol{x})=\sigma$, where $\sigma$ represents the attenuation coefficient in the X-ray scene. The predicted pixel intensity $\hat{I}(\boldsymbol{r})$ of an X-ray image is then described by the Beer-Lambert Law:
@@
where $t_{n}$ and $t_{f}$ represent the near and far thresholds along the ray. The X-ray source is assumed to emit a constant initial intensity $I_{0}$ independent of the viewing direction.
@@ IV-C Static and Dynamic Scene Decomposition
The base of our method is the static and dynamic decoupling of the scene, inspired by the work of D2NeRF and Liu et al. [11, 10]. Specifically, we decouple our scene in a static component $F_{\Theta_{s}}$ and dynamic component $G_{\Theta_{d}}$. Two MLPs represent these with weights $\Theta_{s}$ and $\Theta_{d}$, respectively, as shown in Figure 2 (a).
@@
The static component $F_{\Theta_{s}}$ predicts a static attenuation coefficient $\sigma_{s}$ from a positionally-encoded 3D coordinate $\gamma(\boldsymbol{x})$, similarly to the originally proposed NeRF [5]. We use a coarse-to-fine windowed positional encoding $\gamma$ in our work, which will be further described in Section IV-E. Unlike the static component, the dynamic component $G_{\Theta_{d}}$ is also conditioned on a per-cardiac phase latent code $\tau_{i}$ as input and predicts a dynamic attenuation coefficient $\sigma_{d}$. This latent code $\tau_{i}$ is a learned vector representing a discrete cardiac phase $i$. Specifically, each cardiac cycle, represented by a fixed interval of the ECG as shown in Figure 2 (a), is divided into $T$ equal parts. The parameter $T$, where $i\in\{1,\dots,T\}$, is pre-defined based on the frame rate of the angiogram sequence. In other words, we assume that the time of the frame matches the cardiac phase $i$. Note that, whereas D2NeRF [11] and Liu et al. [10] represent $\tau_{i}$ as per-frame time, we represent $\tau_{i}$ as a cardiac phase in the cardiac cycle. This approach allows us to match cardiac phases across viewpoints.
@@ IV-D Dynamic Reconstruction Components
The function of the static and dynamic decoupling in our work is to separate the scene in the background and the dynamic coronary artery. However, the coronary artery is not the only dynamic element in our CA scene. For example, the heart and ribs also experience motion. Moreover, the sparsity of the structure of the blood vessel also poses a challenge in the separation. Specifically, we propose weighted pixel sampling to address the blood vessel’s separation from the image’s background. We adopt static vs. dynamic factorization loss to correctly separate static and dynamic scenes on a position level [11]. Lastly, to force our dynamic scene to only model the blood vessel, we impose a sparsity structure constraint on our dynamic scene [12]. Figure 2 (c) shows an example of the resulting dynamic reconstruction. As can be seen, the dynamic output mostly consists of the coronary artery structure, while the static output represents the background. Next, we will provide more details on the strategies to achieve these results.
Weighted Pixel Sampling    To optimize our model, we sample rays from all angiogram sequence frames, as was described in Section IV-C. However, the blood vessel structure occupies only a few pixels in every coronary angiogram frame. When all pixels are sampled uniformly, this structure may be missed and, therefore, difficult to separate from the background. We perform weighted pixel sampling to reconstruct the blood vessels rather than the background. The weights are computed based on the assumption that the blood vessel structure is of high contrast and motion. We calculate a sampling probability map based on the variance of each pixel value across the frames of one angiogram sequence. Figure 2 (b) provides an example of a probability map for one angiogram sequence. As shown, the blood vessel structure is distinguished from the background with higher sampling probability, indicated by the darker red areas. During training, we sample a pre-defined fraction of rays $V$ from this map’s thresholded set of high-variance pixels. The remaining pixels are sampled randomly from all possible pixels.
Static vs. Dynamic Factorization Loss    A position in space should be occupied by the static or dynamic scene but never by both. The photometric loss, detailed in Equation 1, does not guarantee this behavior and, therefore, a correct separation of static and dynamic scenes. To enforce this separation, we adopt the Static vs. Dynamic Factorization regularizer from D2NeRF [11]. This regularizer defines a spatial ratio of dynamic and static attenuation coefficients as
@@
D2NeRF also proposes a skewness on top of the binary entropy loss to slightly favor static explanations of the scene, as scenes are naturally more static than dynamic. In our scene, the dynamic component is sparse in structure and occupies only a few points in the 3D space. The skewness measure is generically applied to the 3D space and may lead to a degenerate solution where the dynamic component is completely removed. We, therefore, do not incorporate this skewness measure.
@@
To avoid artifacts from appearing in empty spaces, we disregard non-hitting rays based on their accumulated density, similar to previous work [12]. However, we always compute the entropy for rays likely to hit a blood vessel to avoid these rays nearing zero. We derive this likelihood from the blood vessel probability map, which we also utilize for weighted pixel sampling as described in the respective paragraph and shown in Figure 2 (b).
@@ IV-E Sparse-view Reconstruction Components
The main goal of our work is to reconstruct a coronary artery structure from an extremely sparse amount of views. While the dynamic reconstruction components provide a decoupled reconstruction of the coronary artery, degenerate solutions may still occur when limiting the number of views for training. We, therefore, introduce two components inspired by existing literature for sparse-view reconstruction with NeRFs in natural scenes [8]. Namely, we incorporate windowed positional encoding $\gamma(\textbf{x})$ to achieve a smooth CA scene. Moreover, we also impose edge occlusion regularization on the dynamic scene to avoid common degenerate solutions where artifacts appear near the camera.
Windowed Positional Encoding    A key element in reconstructing high-fidelity scenes with NeRFs is the positional encoding [19]. While these resulting high-frequency inputs allow for fast reconstruction of high-frequency components, they are prone to bias the model to generate high-frequency artifacts, especially in a sparse-view setting [8]. Several works in NeRF have introduced a frequency regularization method, which linearly increases the frequency of the positional encoding over time [42, 43, 8]. This method, which is called windowed positional encoding, forces a coarse-to-fine transition during training. This method reduces high-frequency artifacts [44, 43], also in the case of static and dynamic scene decomposition to enforce smooth separation [11]. Recently, FreeNeRF [8] has shown that their windowed positional encoding is especially effective in avoiding degenerate solutions in an extremely sparse-view setting. We base our windowed positional encoding on the method proposed by FreeNeRF. The positional encoding is then defined as
@@
Dynamic Occlusion Loss    While windowed positional encoding is effective, this strategy still allows for degenerate solutions in the sparse-view setting. Specifically, FreeNeRF demonstrated that NeRF models may generate a solution where high-frequency artifacts appear near the camera when presented with an extremely sparse amount of training views [8]. These solutions are correct for the training views but do not correctly generalize to new viewpoints. To avoid this behavior, FreeNeRF [8] proposes an occlusion regularization loss, punishing high occlusion areas close to the camera. While this loss may be effective in natural scenes with objects in the center, our composite X-ray scene consists of a few non-zero attenuation points. On the other hand, our dynamic scene should be close to empty, with the coronary artery close to the center due to the X-ray system setup. Therefore, we propose to adopt the occlusion regularization loss on our dynamic scene solely. Besides solving the problem of high occlusion areas near the camera, we also expect to achieve solutions where the dynamic scene is nearly empty together with the dynamic entropy loss enforcing sparsity. Figure 3 shows examples of scenes trained with $4$ angiogram sequence, without dynamic occlusion loss $\mathcal{L}_{o}$ and our full $\mathcal{L}_{f}$ method. As can be seen, the occlusion regularization loss forces the dynamic scene to be nearly empty while still modeling the sparse coronary artery structure correctly.
Figure 3: Examples of our method without dynamic occlusion loss and with (full) for $4$ training projections for the static, dynamic, and composite outputs.
@@
where $M(t)$ is a binary mask applied respectively to the distance $t$ of a point along the ray starting from the X-ray source. This mask evaluates to $1$ if the point’s distance $t$ is lower than the distance threshold $D$, otherwise it evaluates to 00. Note that FreeNeRF does not penalize based on a distance $t$ but rather on the first $C$ points starting from the X-ray source. We rather penalize based on a distance threshold $D$, as this is more applicable to a C-arm system setup where physical distances between the X-ray source and patient are measured. Specifically, we expect the points with a distance smaller than $D$ to be outside the patient and, therefore, nearly empty in density, which matches the goal of the occlusion regularization.
IV-F Combined Loss
The loss of our full method $\mathcal{L}_{f}$ for every ray $\boldsymbol{r}$ and cardiac phase $i$ can be summarized as
$\mathcal{L}_{f}(\boldsymbol{r},\boldsymbol{\tau}_{i})=\mathcal{L}_{p}(\boldsymbol{r},\boldsymbol{\tau}_{i})+\lambda_{b}\mathcal{L}_{b}(\boldsymbol{r},\boldsymbol{\tau}_{i})\\ +\lambda_{e}\mathcal{L}_{e}(\boldsymbol{r},\boldsymbol{\tau}_{i})+\lambda_{o}\mathcal{L}_{o}(\boldsymbol{r},\boldsymbol{\tau}_{i}),$
where $\lambda_{b}$, $\lambda_{e}$, and $\lambda_{o}$ represent the weights for the static vs. dynamic factorization, dynamic entropy, and dynamic occlusion loss, respectively. The final model optimization is performed by sampling these rays $\boldsymbol{r}$ from the total set of rays $\mathcal{R}$ defined by the given ground truth pixels at cardiac phases $i$. Rather than utilizing constant weights for $\lambda_{b}$, $\lambda_{e}$, and $\lambda_{o}$, we linearly increase them over time with a pre-determined number of delay steps. This approach is fitting due to the coarse-to-fine transition of the windowed positional encoding, as imposing weights too early in a low-frequency setting may lead to degenerate solutions.

### v2

@@ IV NeRF-CA
Figure 2: Overview of the NeRF-CA method (a). The full loss function $\mathcal{L}_{f}$ (1) consists of (2) Photometric Loss $\mathcal{L}_{p}$, (3) Static vs. Dynamic Factorization Loss $\mathcal{L}_{b}$, (4) Dynamic Entropy Loss $\mathcal{L}_{e}$, and (6) Dynamic Occlusion Loss $\mathcal{L}_{o}$. We also depict the (b) training input and (c) validation output, in yellow and pink, respectively.
In this section, we describe our method, NeRF-CA. Figure 2 (a) shows an overview of our method. We categorize the components of our method according to their function. Specifically, we propose dynamic reconstruction components, shown in green in Figure 2 (a), and sparse-view reconstruction components, shown in purple in Figure 2 (a). Overall, the X-ray scene is expressed as an absorption-only model with attenuation coefficients $\sigma$, as will be further described in Section IV-B. Figure 2 (b, c) depicts the model training input and validation output in yellow and pink, respectively. We will utilize these colors throughout the paper to distinguish the training (yellow) and validation (pink) views.
The basis of our method is decoupling the scene into a static and dynamic scene. The static model represents the background, and the dynamic model represents the dynamic coronary artery structure. The static and dynamic models predict their attenuation coefficients $\sigma_{s}$ and $\sigma_{d}$, respectively. These coefficients are composited to obtain the scene from which the predicted pixel intensities are calculated. Figure 2 (c) depicts the outputs of our methods, specifically, the static, dynamic, and composite outputs for one validation view. Whereas the static model solely relies on the positionally-encoded 3D coordinate as input, the dynamic models also receive a per-cardiac phase latent code $\tau_{i}$. This latent code represents a cardiac phase along the cardiac cycle derived from the simultaneously recorded ECG signal. More details on the scene decomposition can be found in Section IV-C.
By decomposing the scene, we can individually impose regularizers for the static and dynamic scenes, allowing the final dynamic scene to model the coronary artery. In Section IV-A, we provide the overall objective and loss function of our method while introducing these regularizers. Sections IV-D and IV-E describe the regularizers for dynamic and sparse-view reconstruction, respectively. A core aspect of our approach is weighted pixel sampling, which guides the model to distinguish the sparse blood vessel structures from the background. This is achieved using a blood vessel probability map, which estimates the likelihood of each pixel containing a blood vessel. This probability map is shown in Figure 2 (b) and detailed in Section IV-D.
@@ IV-A Overall Objective
Our contribution lies in developing a unique combination of dynamic and sparse-view reconstruction, enabled by an ECG-timed temporal integration of the cardiac sequence and a combination of several regularization techniques targeted to tackle the unique challenges presented by CA data. Figure 2 (c) shows examples of the predicted coronary angiogram views and the ground truth views.
The full loss of NeRF-CA $\mathcal{L}_{f}$ can be summarized as
$\mathcal{L}_{f}(\boldsymbol{r},\boldsymbol{\tau}_{i})=\mathcal{L}_{p}(\boldsymbol{r},\boldsymbol{\tau}_{i})+\lambda_{b}\mathcal{L}_{b}(\boldsymbol{r},\boldsymbol{\tau}_{i})\\ +\lambda_{e}\mathcal{L}_{e}(\boldsymbol{r},\boldsymbol{\tau}_{i})+\lambda_{o}\mathcal{L}_{o}(\boldsymbol{r},\boldsymbol{\tau}_{i}),$ (1)
where $\mathcal{L}_{p}$ represents the photometric loss, $\mathcal{L}_{b}$ represents the static vs. dynamic factorization loss with weight $\lambda_{b}$, $\mathcal{L}_{e}$ represents the dynamic entropy loss with weight $\lambda_{e}$, and $\mathcal{L}_{o}$ represents the dynamic occlusion loss with weight $\lambda_{o}$. Each of these components will be introduced in the next paragraphs. The final model optimization is performed by sampling rays $\boldsymbol{r}$ from a total set of rays $\mathcal{R}$ defined by the ground truth pixels. Each pixel has a cardiac phase $i$ assigned, which is denoted by a time latent code $\boldsymbol{\tau}_{i}$, which will be detailed in Section IV-C.
For dynamic reconstruction, we decouple the static background from the dynamic coronary artery structure, as detailed in Section IV-D. We combine static and dynamic decomposition with a weighted pixel sampling strategy and two regularizers to accurately decouple the coronary artery from the background. Our method is self-supervised through the standard photometric loss $\mathcal{L}_{p}$, computed based on the composition of the static and dynamic components. Weighted pixel sampling is proposed to separate the sparse blood vessel structure from the background at the image level. The static vs. dynamic factorization regularizer $\mathcal{L}_{b}$ is enforced to separate the static and dynamic scene on a 3D point level [14]. In addition, we enforce entropy $\mathcal{L}_{e}$ in the decoupled dynamic scene to achieve the sparse structure of the blood vessel [15].
For sparse-view reconstruction, we reconstruct from an extremely sparse number of angiogram views. The sparse-view reconstruction components, consisting of windowed positional encoding and dynamic occlusion loss, will be detailed in Section IV-E. Windowed positional encoding is a proven strategy to enforce smoothness [9]. We employ this strategy in the decoupled static and dynamic scenes to enforce smoothness in both scenes, avoiding degenerate high-frequency solutions. The occlusion loss $\mathcal{L}_{o}$ punishes high occlusion areas close to the camera, an expected faulty behavior when applying NeRFs for sparse-view reconstruction [9]. We demonstrate that imposing such a loss can benefit our sparse dynamic scene.
@@ IV-B Neural Radiance Field for X-ray
X-ray scenes can be described by the absorption-only model, assuming the Beer-Lambert law [22]. Similarly to other works applying NeRFs for X-ray scenes (e.g., [31]), we approximate the scene by $F_{\Theta}(\boldsymbol{x})=\sigma$, where $\sigma$ represents the attenuation coefficient in the X-ray scene. The predicted pixel intensity $\hat{I}(\boldsymbol{r})$ of an X-ray image is then described as:
@@
where $t$ represents the position along the ray, and $t_{n}$ and $t_{f}$ represent the near and far thresholds along the ray. The X-ray source is assumed to emit a constant initial intensity $I_{0}$ independent of the viewing direction.
@@ IV-C Static and Dynamic Scene Decomposition
The base of our method is the static and dynamic decoupling of the scene, inspired by the work of D2NeRF and Liu et al. [14, 13], and CA reconstruction works that separate the blood vessel from the background during reconstruction [38]. Specifically, we decouple our scene in a static component $F_{\Theta_{s}}$ and dynamic component $G_{\Theta_{d}}$. Two MLPs represent them with weights $\Theta_{s}$ and $\Theta_{d}$, respectively, as shown in Figure 2 (a).
@@
The static component $F_{\Theta_{s}}$ predicts a static attenuation coefficient $\sigma_{s}$ from a positionally-encoded 3D coordinate $\gamma(\boldsymbol{x})$, similar to the originally proposed NeRF [6]. We use a coarse-to-fine windowed positional encoding $\gamma$ in our work, which will be further described in Section IV-E. Unlike the static component, the dynamic component $G_{\Theta_{d}}$ is also conditioned on a latent code per cardiac phase $\tau_{i}$ as input and predicts a dynamic attenuation coefficient $\sigma_{d}$. This latent code $\tau_{i}$ is a learned vector representing a discrete cardiac phase $i$. Specifically, each cardiac cycle, represented by a fixed interval of the ECG as shown in Figure 2 (a), is divided into $T$ equal parts. The parameter $T$, where $i\in\{1,\dots,T\}$ $\bmod$ $T$, is pre-defined based on the frame rate of the angiogram sequence. In other words, we assume that the frame time matches the cardiac phase $i$. A key difference in our approach lies in how we define $\tau_{i}$. While D2NeRF [14] and Liu et al. [13] represent $\tau_{i}$ as per-frame time, we instead represent $\tau_{i}$ as a cardiac phase in the cycle. This approach allows us to match cardiac phases across viewpoints.
@@ IV-D Dynamic Reconstruction Components
The function of the static and dynamic decoupling in our work is to separate the scene in the background and the dynamic coronary artery in a self-supervised manner. However, the coronary artery is not the only dynamic element in our CA scene. For example, the heart and ribs also experience motion. Moreover, the sparsity of the structure of the blood vessel also poses a challenge in the separation. Specifically, we propose weighted pixel sampling to address the blood vessel’s separation from the image’s background. We adopt static vs. dynamic factorization loss to correctly separate static and dynamic scenes on a position level [14]. Lastly, to force our dynamic scene to only model the blood vessel, we impose a sparsity structure constraint on our dynamic scene [15]. An alternative to our self-supervised approach is to directly supervise the dynamic output of our model using a coronary artery segmentation mask, such as one generated by a Frangi filter [39], which we will ablate in Section V-E. Figure 2 (c) shows an example of our dynamic reconstruction. As can be seen, the dynamic output mostly consists of the coronary artery structure, while the static output represents the background. Next, we will provide more details on the strategies to achieve these results.
Weighted Pixel Sampling    To optimize a NeRF model, rays are sampled for every pixel of every frame, as was described in Section IV-C. However, the blood vessel structure occupies only a few pixels in every coronary angiogram frame. When all pixels are sampled uniformly, this structure may be missed and, therefore, difficult to separate from the background. Therefore, we uniquely introduce weighted pixel sampling to reconstruct the blood vessels rather than the background. The weights are computed based on the assumption that the blood vessel structure is of high contrast and motion. We calculate a sampling probability map based on the variance of each pixel value across the frames of one angiogram sequence. Figure 2 (b) provides an example of a probability map for one angiogram sequence. The blood vessel structure is distinguished from the background with higher sampling probability, indicated by the darker red areas. During training, we sample a pre-defined fraction of rays $V$ from this map’s thresholded set of high-variance pixels. The remaining pixels are sampled randomly from all possible pixels.
Static vs. Dynamic Factorization Loss    A position in space should be occupied by the static or dynamic scene but never by both. The photometric loss, detailed in Equation 2, does not guarantee this behavior and, therefore, a correct separation of static and dynamic scenes. To enforce this separation, we adopt the Static vs. Dynamic Factorization regularizer [14]. This regularizer defines a spatial ratio of dynamic and static attenuation coefficients as
@@
D2NeRF uses a skewness term on top of the binary entropy loss to favor static explanations, assuming scenes are mostly static. However, since our dynamic component is sparse and occupies only a few points in the 3D space, applying skewness leads to degenerate solutions where the dynamic component is entirely removed. We omit the skewness term and instead achieve structure sparsity uniquely through a dynamic entropy loss, as detailed below.
@@
To avoid artifacts from appearing in empty spaces, we disregard non-hitting rays based on their accumulated density, similar to previous work [15]. However, given that we want to achieve sparseness for the areas that have blood vessels, we always compute entropy for rays likely to hit a blood vessel, avoiding these rays to near zero. We derive this likelihood from the blood vessel probability map, which we also utilize for weighted pixel sampling as described in the respective paragraph and shown in Figure 2 (b).
@@ IV-E Sparse-view Reconstruction Components
The main goal of our work is to reconstruct a coronary artery structure from an extremely sparse amount of views. While the dynamic reconstruction components provide a decoupled reconstruction of the coronary artery, degenerate solutions may still occur when limiting the number of views for training. We, therefore, introduce two components for sparse-view reconstruction. Namely, we incorporate windowed positional encoding $\gamma(\textbf{x})$ to achieve a smooth CA scene. Moreover, we also impose edge occlusion regularization on the dynamic scene to avoid common degenerate solutions where artifacts appear near the camera.
Windowed Positional Encoding    A key element in reconstructing high-fidelity scenes with NeRFs is the positional encoding [21]. While these resulting high-frequency inputs allow for fast reconstruction of high-frequency components, they are prone to bias the model to generate high-frequency artifacts, especially in a sparse-view setting [9]. Several works in NeRF have introduced a frequency regularization method, which linearly increases the frequency of the positional encoding over time [41, 42, 9], enforcing smoothness in the scene. Recently, FreeNeRF [9] has shown that their windowed positional encoding is especially effective in avoiding degenerate solutions in an extremely sparse-view setting. This windowed positional encoding is applied on both the static and dynamic component of our scene, which enforces smoothness in each decoupled component. The positional encoding is defined as
@@
Dynamic Occlusion Loss    While windowed positional encoding is effective, this strategy still allows for degenerate solutions in the sparse-view setting. Specifically, FreeNeRF [9] demonstrated that NeRF models may generate a solution where high-frequency artifacts appear near the camera when presented with an extremely sparse amount of training views. These solutions are correct for the training views but do not correctly generalize to new viewpoints. To avoid this behavior, FreeNeRF [9] proposes an occlusion regularization loss, punishing high occlusion areas close to the camera. While this loss may be effective in natural scenes with objects in the center, our composite X-ray scene consists of a few non-zero attenuation points. As a result, applying this loss directly to the composite scene would not avoid degenerate outcomes. In contrast, our learned dynamic scene, representing the coronary artery, is expected to be nearly empty, with the coronary artery positioned near the center due to the X-ray system setup. Therefore, on top of our self-supervised decomposition approach, we apply occlusion regularization exclusively to the learned dynamic scene.
@@
where $M(t)$ is a binary mask applied respectively to the distance $t$ of a point along the ray starting from the X-ray source. This mask evaluates to $1$ if the point’s distance $t$ is lower than the distance threshold $D$, otherwise it evaluates to 00. Note that FreeNeRF [9] does not penalize based on a distance $t$ but rather on the first $C$ points starting from the X-ray source. We rather penalize based on a distance threshold $D$, as this is more applicable to a C-arm system setup where physical distances between the X-ray source and patient are known. Specifically, we enforce the points with a distance smaller than $D$ to be outside the patient and, therefore, nearly empty in density, which matches the goal of the occlusion regularization.
Equation 1 provides an overview of NeRF-CA’s full loss function. Rather than utilizing constant weights for the hyperparameters $\lambda_{b}$, $\lambda_{e}$, and $\lambda_{o}$, we linearly increase them over time with a pre-determined number of delay steps. This approach is fitting due to the coarse-to-fine transition of the windowed positional encoding, as imposing weights too early in a low-frequency setting may lead to degenerate solutions.